# SPER: Accelerating Progressive Entity Resolution via Stochastic Bipartite Maximization

### v1

@@ 1. Introduction
To address the high initialization cost of Progressive ER, we propose SPER, a high-velocity framework that integrates semantic embeddings with a continuous stochastic bipartite maximization strategy to prioritize candidate pairs. By fundamentally redefining prioritization as a weighted sampling problem rather than a global ranking one, our approach shifts the paradigm from deterministic ordering to probabilistic filtering, which effectively enables the immediate, linear-time identification of high-confidence matches without any initialization latency.
@@
  • We introduce a novel paradigm for Progressive ER, which replaces the deterministic sorting during the initialization phase with a stochastic process that efficiently retrieves the top-weighted candidate pairs from a set of blocks.
  • We theoretically prove that our stochastic relaxation effectively emulates the optimal sorting order in linear time, due to its constant time complexity per query. As a result, it significantly reduces the prioritization complexity compared to the super-linear costs of existing progressive ER methods.
@@
The remainder of this paper is organized as follows. Section 2 reviews the major existing work in Progressive ER. Section 3 formally defines the problem of scalable utility maximization on bipartite graphs. Section 4 introduces the SPER framework, detailing the stochastic bipartite maximization strategy and providing the theoretical proof of its convergence to the optimal sorting order. Section 5 presents the comprehensive experimental evaluation, comparing SPER against state-of-the-art baselines on eight real-world datasets. Finally, Section 6 concludes the paper and outlines directions for future research.
@@ 3. Problem Formulation
  • $\mathcal{W}$ denotes the set of edge weights, with each $w_{(r,s)}\in[0,1]$ denoting the matching likelihood between profiles $r\inRR$ and $s\inSS$.
@@ 4. Approach
To this end, SPER embeds the records of $R$ into dense vectors using an embedding model and then, it stores these embeddings in an Approximate Nearest Neighbor Search (ANNS) index capable of returning the top-$k$ neighbors in logarithmic query time. More specifically, it SPER involves three phases:
@@
    Latency Indeterminacy: Strictly selecting the top-$k$ pairs requires observing the entire candidate set to establish a ranking, forcing a batch-processing model with $\mathcal{O}(n\log n)$ sorting costs, where $n$ is the total number of candidate pairs (i.e., $n=k\cdot|S|$). Stochastic sampling approximates this ranking in $\mathcal{O}(n)$ time.
@@ 4.1. Stochastic Bipartite Maximization
We first establish the target budget $B$ as a fixed fraction $\rho$ of the total expected candidate volume, such that $B=\rho\cdot k|S|$. Ideally, the scaling factor $\alpha$ would be set to satisfy the constraint $\sum p_{(r,s)}=B$ exactly; however, computing this optimum is time-consuming for large-scale input datasets. Instead, SPER initializes $\alpha$ using a conservative estimate derived directly from the budget definition: $\alpha\approx B/(0.5\cdot k|S|)=2\rho$, where $0.5$ serves as a safety prior for the average similarity weight. This initialization ensures that our approach begins in a state of controlled under-utilization and ramps up, avoiding an initial overflow that would require drastic correction. Subsequently, to maintain this budget dynamically, SPER employs an online adaptive calibration where candidate pairs are processed in windows of size $W$. After each window, we compare the observed number of selections $m_{w}$ to a small fixed target $B_{w}=\left\lceil B\cdot\frac{W}{|S|}\right\rceil$ and update $\alpha$ multiplicatively:
(3) $\alpha=\alpha\Bigl(1+\eta\frac{B_{\mathrm{w}}-m_{\mathrm{w}}}{B_{\mathrm{w}}}\Bigr),$
@@
Having fixed the selection probabilities, let $m=|\mathcal{S}^{\prime}|$ denote the random number of candidate pairs selected by the algorithm. Modeling the selection process as a sequence of independent Bernoulli trials [^2] Given that the probabilities vary for each candidate pair, the random variable $m$ follows a Poisson Binomial Distribution., each with parameter $p_{(r,s)}$, its expectation and variance are:
@@
For concentration, Chernoff bounds for sums of independent Bernoulli variables (with mean $\mu=\mathbb{E}[m]=B$) yield, for $0<\epsilon\leq 1$,
@@
1:Input: Dataset $S$, index $I$, neighbors $k$, window size $W$, selection percentage $\rho$, embedding model $\mathcal{T}$
@@
18:   $\alpha\leftarrow\alpha(1+\eta\frac{B_{w}-m_{w}}{B_{w}})$ $\triangleright$ Update scaling factor $\alpha$
@@ 4.2. The SPER Algorithm
    Retrieval (Lines 57): For each query entity $s$ in $S$, SPER embeds it into a dense vector using embedding model $\mathcal{T}$ and queries the index II to retrieve the set $\mathcal{C}_{s}$ of the Id’s of its top-$k$ nearest candidates of $R$.
@@
The total runtime is composed of two phases: (1) the retrieval phase, where for every entity in $S$, querying the index takes logarithmic time with respect to the index size $|R|$. Across all queries, this sums to $\mathcal{O}(|S|\cdot\log|R|)$, and (2) the selection phase, where for each of the $|S|$ queries, the algorithm processes $k$ candidates. The stochastic check at Lines 10 and 11 is a constant time operation $\mathcal{O}(1)$. The total selection effort is thus $\mathcal{O}(k\cdot|S|)$. Therefore, the combined time complexity is:
(7) $T_{\text{total}}=\mathcal{O}(|S|\cdot\log|R|+k\cdot|S|).$
@@ 5.1. Experimental results
Figure 4 demonstrates that SPER and the baselines deliver almost comparable recall for the smallest budgets across all benchmarks. We report $B$ in absolute terms, representing the direct output determined by the corresponding relative factor $\rho$. A steeper curve indicates superior progressiveness, as it signifies that a higher percentage of true matches is identified earlier in the emission process. However, as the budget increases, SPER consistently outperforms its competitors on the more complex datasets, showing significant average improvements on Abt-Buy ($12\%$), Amazon-Google ($9\%$), DBLP-ACM ($13\%$), DBLP-Scholar ($10\%$), and Walmart-Amazon ($7\%$). These performance gains stem from SPER’s use of semantic embeddings, which allow it to link entities that are lexically distinct but semantically identical (e.g., entities with PVLDB vs. Proceedings of the VLDB Endowment). In contrast, the three baselines rely on the token overlap, which limits their recall on these semantically heterogeneous datasets. Conversely, on NC-Voters and DBLP, where the differences between matching entities arise primarily from synthetic lexical perturbations, the baselines slightly outperform SPER by $5\%$ and $1\%$, on average, respectively.

### v2

@@ 1. Introduction
To address the high initialization cost of Progressive ER, we propose SPER, a high-velocity framework that integrates semantic embeddings with a continuous stochastic bipartite maximization strategy to prioritize candidate pairs. By fundamentally redefining prioritization as a weighted sampling problem rather than a global ranking one, our approach shifts the paradigm from deterministic ordering to probabilistic filtering. This enables the immediate, linear-time identification of high-confidence matches with utility statistically equivalent to the optimal ranking, yet without any initialization latency. Even though SPER might not pick the exact same list of pairs as global sorting, there is no significant impact on effectiveness, as the total similarity weight of the selected pairs is equally high. We experimentally demonstrate the insignificant (if any) impact on effectiveness along with the significant gains in time efficiency through an extensive experimental study that involves 8 datasets commonly used in the literature.
@@
  • We introduce a novel paradigm for Progressive ER, which replaces the deterministic sorting during the initialization phase with a stochastic process that efficiently retrieves the top-weighted candidate pairs by dynamically scaling selection probabilities to adhere to a strict budget constraint.
  • We theoretically prove that our stochastic relaxation efficiently retrieves high-weight candidates in linear time, bypassing the expensive sorting bottleneck while maintaining theoretical utility guarantees. As a result, it significantly reduces the prioritization complexity compared to the super-linear costs of existing progressive ER methods.
@@
The remainder of this paper is organized as follows. Section 2 reviews the major existing work in Progressive ER. Section 3 formally defines the problem of scalable utility maximization on bipartite graphs. Section 4 introduces the SPER framework, detailing the stochastic bipartite maximization strategy and providing the theoretical proof of its convergence to the expected utility of the optimal baseline. Section 5 presents the comprehensive experimental evaluation, comparing SPER against state-of-the-art baselines on eight real-world datasets. Finally, Section 6 summarizes our key findings and concludes the work.
@@ 3. Problem Formulation
  • $\mathcal{W}$ is the set of edge weights, where each $w_{(r,s)}\in[0,1]$ indicates the matching likelihood between profiles $r\in R$ and $s\in S$.
@@ 4. Approach
To this end, SPER embeds the records of $R$ into dense vectors using an embedding model and then, it stores these embeddings in an Approximate Nearest Neighbor Search (ANNS) index capable of returning the top-$k$ neighbors in logarithmic query time. More specifically, SPER involves three phases:
@@
    Latency Indeterminacy: Strictly selecting the top-$B$ pairs requires observing the entire candidate set to establish a ranking, forcing a batch-processing model with $\mathcal{O}(n\log n)$ sorting costs, where $n$ is the total number of candidate pairs (i.e., $n=k\cdot|S|$). Stochastic sampling approximates the utility of this optimal selection in $\mathcal{O}(n)$ time.
@@ 4.1. Stochastic Bipartite Maximization
We first establish the target budget $B$ as a fixed fraction $\rho$ of the total expected candidate volume, such that $B=\rho\cdot k|S|$. Ideally, the scaling factor $\alpha$ would be set to satisfy the constraint $\sum p_{(r,s)}=B$ exactly; however, computing this optimum is time-consuming for large-scale input datasets.
Instead, SPER initializes $\alpha$ using a conservative estimate derived directly from the budget definition: $\alpha\approx B/(0.5\cdot k|S|)=2\rho$, where $0.5$ serves as a safety prior for the average similarity weight. This initialization ensures that our approach begins in a state of controlled under-utilization and ramps up, avoiding an initial overflow that would require drastic correction. Subsequently, to maintain this budget dynamically, SPER employs an online adaptive calibration where candidate pairs are processed in windows of size $W$. After each window, we compare the observed number of selections $m_{w}$ to a small fixed target $B_{w}=\left\lceil B\cdot\frac{W}{|S|}\right\rceil$ and update $\alpha$ multiplicatively:
(3) $\alpha_{\text{new}}=\alpha_{\text{old}}\Bigl(1+\eta\frac{B_{\mathrm{w}}-m_{\mathrm{w}}}{B_{\mathrm{w}}}\Bigr),$
@@
Having fixed the selection probabilities, let $m$ denote the random number of candidate pairs selected by the algorithm. Modeling the selection process as a sequence of independent Bernoulli trials [^2] Given that the probabilities vary for each candidate pair, the random variable $m$ follows a Poisson Binomial Distribution., each with parameter $p_{(r,s)}$, its expectation and variance are:
@@
To rigorously quantify this stability, Chernoff bounds for sums of independent Bernoulli variables (with mean $\mu=\mathbb{E}[m]=B$) yield, for $0<\epsilon\leq 1$,
@@
1:Input: Dataset $S$, indexed dataset $R$ as $I$, neighbors $k$, window size $W$, selection percentage $\rho$, embedding model $\mathcal{T}$
@@
18:   $\alpha_{\text{new}}\leftarrow\alpha_{\text{old}}(1+\eta\frac{B_{w}-m_{w}}{B_{w}})$ $\triangleright$ Update scaling factor $\alpha$
@@ 4.2. The SPER Algorithm
    Retrieval (Lines 57): For each query entity $s$ in $S$, SPER embeds it into a dense vector using embedding model $\mathcal{T}$ and queries the index II to retrieve the set $\mathcal{C}_{s}$ of the Id’s of its top-$k$ nearest candidates of RR. [^4] Embedding and retrieval are batched operations.
@@
The total runtime is composed of two phases: (1) the retrieval phase, where for every entity in $S$, querying the index takes logarithmic time with respect to the index size $|R|$. Across all queries, this sums to $\mathcal{O}(|S|\cdot\log|R|)$, and (2) the selection phase, where for each of the $|S|$ queries, the algorithm processes $k$ candidates. The stochastic check at Lines 10 and 11 is a constant time operation $\mathcal{O}(1)$. The total selection effort is thus $\mathcal{O}(k\cdot|S|)$. Therefore, the combined time complexity is $\mathcal{O}(|S|\cdot\log|R|+k\cdot|S|)$.
@@ 5.1. Experimental results
Figure 4 demonstrates that SPER and the baselines deliver almost comparable recall for the smallest budgets across all benchmarks. We report $B$ in absolute terms, representing the direct output determined by the corresponding relative factor $\rho$. A steeper curve indicates superior progressiveness, as it signifies that a higher percentage of true matches is identified earlier in the emission process. However, as the budget increases, SPER consistently outperforms its competitors on the more complex datasets, showing significant average improvements on Abt-Buy ($12\%$), Amazon-Google ($9\%$), DBLP-ACM ($13\%$), DBLP-Scholar ($10\%$), and Walmart-Amazon ($7\%$). These performance gains stem from SPER’s use of semantic embeddings, which allow it to link entities that are lexically distinct but semantically identical (e.g., entities with PVLDB vs. Proceedings of the VLDB Endowment). In contrast, the three baselines rely on the token overlap, which limits their recall on these semantically heterogeneous datasets. Conversely, on NC-Voters and DBLP, where the differences between matching entities arise primarily from synthetic lexical perturbations, the baselines slightly outperform SPER by $5\%$ and $1\%$, on average, respectively. We also plot the sorted baseline using embeddings, which prioritizes candidates strictly by their semantic similarity scores in descending order. This serves as an empirical benchmark, demonstrating that SPER’s probabilistic sampling sacrifices negligible effectiveness compared to the computationally expensive deterministic optimal approach.